You are here: Foswiki>LabstaffWeb Web>LabstaffDowntime>DowntimeHistory>Outage201104 (02 May 2011, lindss2@LAB.CS.RPI.EDU)EditAttach

April 2011 Outages


2011/04/30 OK Home directories unavailable, synchronizing disks
2011/04/13 OK A networking switch in Amos Eaton croaked this morning. Servers have been relocated onto another functional unit
2011/04/11 OK mail server disk problems
2011/04/01 OK home directory server hung, may be due to load from 2 remote access machines. Investigating
filesystem check being performed, server will be back online shortly
2011/04/01 - 6:40am OK mailserver0 reboot to clear a failing disk channel
2011/03/31 OK mailserver1 unresponsive -- bios logs filled

-- StevenLindsey - 2011-05-02
Topic revision: r1 - 02 May 2011, lindss2@LAB.CS.RPI.EDU
 

This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback